OpenStudy (sidney18):

A ladder is leaning against a wall. The top of the ladder is 9 feet (ft) above the ground. If the bottom of the ladder is moved 3ft farther from the wall, the ladder will be lying flat on the ground, still touching the wall. How long, in feet, is the ladder?

OpenStudy (welshfella):



OpenStudy (welshfella):



OpenStudy (welshfella):

Do you see why one leg of the right angled triangle = x - 3 ft?

OpenStudy (sidney18):

yes. To find that leg you have to subtract the 3ft from what ever x (the length of the segment) is to find the length of that leg.

OpenStudy (welshfella):

yes - x is the length of the ladder now we apply the pythagoras theorem to the triangle x^2 = (x - 3)^2 + 9^2 solve this for x and you have the answer

OpenStudy (sidney18):

The ladder is 15 ft tall?

OpenStudy (welshfella):

right